The measure of an angle is 15 less than twice the measure of its complement. Find the measure of each angle.
Answer by cleomenius(959)

2(90 - x) - 15 = x
180 -2x - 15 = x
165 = 3x
x = 55 the angle.
35 the complement.
2x = 70
70 - 15 = 55 degrees (check)
Cleomenius.